# PF3plat: Pose-Free Feed-Forward 3D Gaussian Splatting

### v1

@@ 4.4 Ablation Study
(I) PFSplat 22.347 0.763 0.205 1.965 0.751 10.113 4.785
(II) - Depth Refinement 21.963 0.759 0.208 2.240 1.134 9.701 3.044
(III) - Pose Refinement 21.519 0.737 0.222 2.349 1.175 12.123 6.347
(IV) - Geometry Confidence 21.239 0.737 0.223 2.303 0.922 9.179 3.533
@@
(I-III) - Tri. Consis. Loss 18.832 0.6247 0.418 5.787 2.105 17.117 9.422
(I-IV) - Regularization Loss 20.981 0.722 0.243 4.698 1.908 11.172 8.384
@@
Table 4: Component ablations on RealEstate10K.
From the results, we find that our method improves by a large margin when comparing (0) with (I). This improvement is further supported by the comparisons from (I) to (IV) and from (I-III) to (I-IV), which show performance degradation as each component is removed. We also demonstrate that without pre-trained weights for the depth and correspondence networks, the training either fails or achieves significantly lower performance. Similar observations are made in (I-I), (I-II), and (I-V), where we identify that directly tuning the depth network or training only with photometric losses leads to failure in the training process. The former issue may arise from overfitting, a common problem when directly manipulating foundation models. With only the photometric loss, we observe that after certain iterations, as the baseline becomes wider, the training loss quickly diverges.
Figure 3: Qualitative results on RealEstate10K and ACID dataset. Given two context views (a) and (b), we compare novel view rendering results with other methods. We provide more qualitative results on RealEstate10K, ACID and DL3DV in supplementary material.
@@
Splatt3R 15.636 0.502 0.360 1.312 0.521 8.715 1.891 20
Ours 22.347 0.763 0.205 1.965 0.751 10.113 4.785 0.390
Ours + TTO 23.132 0.779 0.202 1.965 0.814 9.996 4.701 13
(a) Performance and speed comparisons on RealEstate-10K against DUSt3R variants.
@@
Ours 26.3055 0.8664 0.1230 0.0011624 27.0822 0.8895 0.1055 0.0005684
(c) RealEstate10K 6, 12 input views.
@@
Ours 20.542 0.647 0.267 0.0672 0.0263 9.5373 5.3940 21.086 0.708 0.234 0.052 0.027 13.133 8.095
(d) Cross-dataset Evaluation.
@@ 4.5 Analysis and More Results
Comparison to DUSt3R extensions.
In this study, we compare our results with those of InstantSplat (Fan et al., 2024) and Splatt3R (Smart et al., 2024). Although these are preprints, we include this comparison for completeness, as their tasks and methods are closely related to ours. The results are summarized in Tab. 5(a). We find that InstantSplat achieves superior performance in novel view synthesis compared to our method; however, when we adopt a similar test-time optimization (TTO) strategy, our approach requires significantly less optimization time to achieve comparable or better results. Additionally, our method infers substantially faster without TTO, demonstrating its high practicality for real-time applications. Splatt3R, a concurrent work built on a similar concept to ours, requires ground-truth depth and pose for training, which are not available in the RealEstate10K dataset. Due to scale discrepancies, their performance is significantly lower, and without a mechanism to handle the scale differences, an issue not addressed in the original paper, the results are not directly comparable to ours.
@@
We conduct a comprehensive inference speed comparison between our method and competing approaches using varying numbers of input images, specifically $N\in{2,6,12}$. For each scenario, we evaluate the time required to render 1, 3, and 5 views. The results, summarized in Tab. 5(b), show that our approach is generally faster than existing methods. However, for $N=12$, our inference speed is slower than that of DBARF, as our method involves estimating camera poses via a robust solver for every pairwise combination. Despite this overhead, our approach gains a significant advantage as the number of rendered views increases, due to the efficient rendering capabilities of 3DGS once the scene has been fully reconstructed. Finally, in order to provide a detailed breakdown, we measured three values for our method: overall inference time, UniDepth processing time, and decoder time, isolating the contributions of each component to the total runtime. Given 2, 6 and 12 views and to render a single target view, it takes 0.251, 0.832 and 1.535 seconds for UniDepth inference, while it takes approximately consistent 0.00247 seconds for rendering.
@@
In practical scenarios, more than two views ($N>2$) are commonly used. Therefore, we demonstrate that our method can process multiple views and render $I_{t}$. We input $N$ views into our network to obtain $\hat{\mathcal{P}}_{i}$, $\hat{\mathcal{D}}_{i}$, and $(\mu_{i},\sigma_{i},\Sigma_{i},c_{i})$. Following a similar approach to Chen & Lee (2023), we select the top-$k$ nearby views using $\hat{\mathcal{P}}_{i}$ and render $\hat{I}_{t}$ to compare with the ground truth target view image. For this evaluation, we compare our method with those of Chen & Lee (2023) and Smith et al. (2023), since the method by Hong et al. (2024) can only take two input views. We also report the Absolute Trajectory Error (ATE). The results are summarized in Tab. 5(c). From these results, we find that our method achieves significantly better performance than the others, highlighting our capability to extend to multiple $N$ views.
@@
To demonstrate the generalization capability of our method, we conduct a cross-dataset evaluation and compare against Hong et al. (2024). Specifically, we evaluate performance on RealEstate10K and DL3DV, using each dataset for training in a cross-dataset setting. The results, summarized in Tab. 5(d), show that our method achieves a PSNR of over 20 dB for both datasets, significantly outperforming Hong et al. (2024). This indicates that, even under out-of-distribution conditions, our method produces high-quality renderings, highlighting its robustness and effectiveness in zero-shot capability.
@@ 5 Conclusion
In this paper, we have introduced learning-based framework that tackles pose-free novel view synthesis with 3DGS, enabling efficient, fast and photorealistic view synthesis from unposed images. The proposed framework, PFSplat, is built on the base model comprising of foundation models to overcome inherent limitations of 3DGS. While the devised base model already surpasses the existing methods, we have also devised modules to address the limitations of the base model, enhancing the overall performance. This method is capable of training and inference solely from unposed images, even in scenarios where only a handful of images with minimal overlaps are given. We have shown that our approach surpasses all existing methods on real-world large-scale datasets, establishing new state-of-the-art performance.
@@ A.1 Training Details
A.2 More Qualitative Results
Fig. 4, Fig. 5 and Fig. 6 present more novel view rendering results of different methods. On these datasets, our method yields outcomes that are sharper and more geometrically accurate.
A.3 Limitations and Future Work
As our model currently lacks a mechanism to handle dynamic scenes, it is unable to accurately capture scene dynamics or perform view extrapolation. Additionally, our model’s performance is contingent on the quality of the coarse alignments, which rely on the accuracy of the depth and correspondence models. In cases where either of these models fails, our approach may not function optimally. However, because our refinement modules are lightweight, simple, and model-agnostic, incorporating more advanced methods for coarse alignment could further enhance performance.

### v2

@@ 4.4 Ablation Study
(I) PFSplat 23.589 0.782 0.181 1.756 0.897 9.474 4.628
(II) - Depth Refinement 22.012 0.754 0.203 2.342 1.122 9.881 4.952
(III) - Pose Refinement 21.623 0.744 0.219 2.310 1.233 11.889 6.544
(IV) - Geometry Confidence 21.443 0.741 0.223 2.228 1.001 11.322 5.998
@@
(I-III) - Tri. Consis. Loss 19.001 0.644 0.402 5.661 2.099 18.332 10.331
(I-IV) - Regularization Loss 21.332 0.733 0.231 4.555 2.012 12.338 9.998
@@
(a) Performance and speed comparisons on RealEstate-10K against per-scene optimization methods.
@@
CF-3DGS 14.024 0.455 0.450 13.278 8.486 106.397 106.337 25
Ours 23.589 0.782 0.181 1.756 0.897 9.474 4.628 0.390
Ours + TTO 24.689 0.798 0.167 1.662 0.871 8.998 4.311 24
@@
Ours 27.0284 0.8788 0.1158 0.0010048 28.1334 0.9934 0.0988 0.0004228
(d) Cross-dataset Evaluation.
@@
Ours 21.332 0.678 0.234 3.248 1.313 9.432 5.112 21.877 0.733 0.221 2.778 1.511 12.881 7.882
From the results, we find that significant improvement from (0). This improvement is further supported by the comparisons from (I) to (IV) and from (I-III) to (I-IV), which show performance degradation as each component is removed. While we observe that (0) surpasses our final model in terms of the median rotation and translation angular difference, this aligns with a commonly observed trend between learning-based and robust solver-based approaches (Rockwell et al., 2024). Classical methods tend to achieve higher precision, whereas learning-based approaches generally offer greater robustness.
We also demonstrate that without pre-trained weights for the depth and correspondence networks, the training either fails or achieves significantly lower performance. Similar observations are made in (I-I), (I-II), and (I-V), where we identify that directly tuning the depth network or training only with photometric losses leads to failure in the training process. The former issue may arise from overfitting, a common problem when directly manipulating foundation models. With only the photometric loss, we observe that after certain iterations, as the baseline becomes wider, the training loss quickly diverges.
@@ 4.5 Analysis and More Results
Comparison to scene-specific optimization methods.
In this study, we also compare our method with InstantSplat (Fan et al., 2024) and CF-3DGS (Fu et al., 2023), as summarized in Tab. 5(a). Our approach already surpasses InstantSplat, a method that adopts similar 2-stage approach as ours, but instead of feed-forward inference, it iteratively optimizes the 3D Gaussian parameters. This results highlights the effectiveness of our refinement modules and our design. The performance gap widens further when we adopt a similar test-time optimization (TTO) strategy. By using our predictions as initialization, TTO takes significantly less time than InstantSplat, demonstrating high practicality. Finally, CF-3DGS struggles to find accurate camera poses and suffers in rendering quality, likely because its design does not adequately handle wide-baseline images.
@@
We conduct a comprehensive inference speed comparison between our method and competing approaches using varying numbers of input images, specifically $N\in{2,6,12}$. For each scenario, we evaluate the time required to render 1, 3, and 5 views. The results, summarized in Tab. 5(b), show that our approach is generally faster than existing methods. However, for $N=12$, our inference speed is slower than that of DBARF, as our method involves estimating camera poses via a robust solver for every pairwise combination. Despite this overhead, our approach gains a significant advantage as the number of rendered views increases, due to the efficient rendering capabilities of 3DGS once the scene has been fully reconstructed. Finally, we provide the inference time of each of our components: overall inference time, UniDepth processing time, and decoder time. Given 2, 6, 12 views and to render a single target view, it takes 0.251, 0.832, 1.535 seconds for UniDepth inference, while rendering takes approximately consistent 0.00247 seconds.
@@
In practical scenarios, more than two views ($N>2$) are commonly used. Moreover, unlike (Ye et al., 2024), our method naturally supports extension to $N$-views, without needing to train our network again. Here, we demonstrate that our method can process multiple views and render $I_{t}$. We input $N$ views into our network to obtain $\hat{\mathcal{P}}_{i}$, $\hat{\mathcal{D}}_{i}$, and $(\mu_{i},\sigma_{i},\Sigma_{i},c_{i})$. Following a similar approach to (Chen & Lee, 2023), we select the top-$k$ nearby views using $\hat{\mathcal{P}}_{i}$ and render $\hat{I}_{t}$ to compare with the ground truth target view image. For this evaluation, we compare our method with those of (Chen & Lee, 2023) and (Smith et al., 2023), since the method by (Hong et al., 2024) can only take two input views. We also report the Absolute Trajectory Error (ATE). The results are summarized in Tab. 5(c). From these results, we find that our method achieves significantly better performance than the others, highlighting our capability to extend to multiple $N$ views. We also provide qualitative results and videos for $N$-view experiments, in supplementary material.
@@
To demonstrate the generalization capability, we conduct a cross-dataset evaluation and compare against (Hong et al., 2024). Specifically, we evaluate performance on RealEstate10K and DL3DV, using each dataset for training in a cross-dataset setting. The results, summarized in Tab. 5(d), show that our method achieves a PSNR of over 20 dB for both datasets, significantly outperforming (Hong et al., 2024). This indicates that, even under out-of-distribution conditions, our method produces high-quality renderings, highlighting its robustness and effectiveness in zero-shot capability.
@@ 5 Conclusion
In this paper, we have introduced learning-based framework that tackles pose-free novel view synthesis with 3DGS, enabling efficient, fast and photorealistic view synthesis from unposed images. Our framework, PFSplat, is built on foundation models to overcome inherent limitations of 3DGS. We have demonstrated that PF3plat is capable of training and inference solely from unposed images, even in scenarios where only a handful of images with minimal overlaps are given. Furthermore, PF3plat surpasses all existing methods on real-world large-scale datasets, establishing new state-of-the-art performance.
Acknowledgement
This research was supported by Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2019-II190075, RS-2024-00509279, RS-2025-II212068, RS-2023-00227592, RS-2024-00457882) and the Culture, Sports, and Tourism R&D Program through the Korea Creative Content Agency grant funded by the Ministry of Culture, Sports and Tourism (RS-2024-00345025, RS-2024-00333068), and National Research Foundation of Korea (RS-2024-00346597).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning, which focuses in practical novel view synthesis. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ A.1 Training Details
A.2 Implementation Details
We provide a detailed illustration in Fig. 2 for our fine alignment and confidence estimation modules.
A.3 Evaluation Details
For the evaluation on RealEstate-10K and ACID, we follow the protocol outlined by (Hong et al., 2024), where evaluation is conducted using unposed triplet images $(I_{1},I_{2},I_{t})$. The test set is divided into three groups, small, middle, and large, based on the extent of overlap between $I_{1}$ and $I_{2}$. This allows the model’s performance to be assessed under varying levels of difficulty, reflecting different real-world scenarios. For the relatively new DL3DV dataset, we introduce a new evaluation protocol. For each scene, we select two context images, $I_{1}$ and $I_{2}$, by skipping frames at intervals of 5 and 10, creating two groups per scene, each representing small and large overlap cases. We then randomly select three target images from the sequence between the context images.
For evaluation metrics, we use standard image quality measures, PSNR, SSIM, LPIPS, and MSE, for novel view synthesis. For camera pose estimation, we compute the geodesic rotation error and angular difference in translation, as commonly done in classical methods (Nistér, 2004; Melekhov et al., 2017). Our statistical analysis reports both the average and median errors, with the median providing robustness against outliers.
A.4 Additional Experiment
Different Strategy for Coarse Alignment.
In this additional experiment, we analyze the impact of different coarse alignment strategies. Specifically, we replace LightGlue with RoMa (Edstedt et al., 2024) and UniDepth with DepthPro (Bochkovskii et al., 2024) for coarse alignment. The results, presented in Tab. 6, reveal that using more robust and accurate correspondences leads to noticeable improvements, as these correspondences enhance our loss computation. However, we observed that integrating DepthPro results in significant performance degradation. This issue may stem from the feature maps used by our fine alignment modules, which could potentially be mitigated through further engineering refinements.
A.5 More Qualitative Results
Fig. 4, Fig. 5 and Fig. 6 present more novel view rendering results of different methods. Fig. 7 shows qualitative results on pose estimation in $N$-view experiments, compared against FlowCAM (Smith et al., 2023). Finally, we provide 6-view and 12-view qualitative results in Fig. 8 and Fig. 9, while videos can be found in the attached file.
Method PSNR SSIM LPIPS Rot. ($\downarrow$) Trans. ($\downarrow$)
Avg. Med. Avg. Med.
MASt3R - - - 2.555 0.751 9.775 2.830
RoMa - - - 2.470 0.391 8.047 1.601
Ours 23.589 0.782 0.181 1.756 0.897 9.474 4.628
DepthPro + Fine Align. (Ours) 21.882 0.701 0.289 2.778 0.812 11.322 5.414
RoMa + Fine Align. (Ours) 24.412 0.799 0.167 2.152 0.501 7.544 3.233
Table 6: Different Strategies for Coarse Alignment.
A.6 Limitations and Future Work
As our model currently lacks a mechanism to handle dynamic scenes, it is unable to accurately capture scene dynamics or perform view extrapolation. Additionally, our model’s performance is contingent on the quality of the coarse alignments, which rely on the accuracy of the depth and correspondence models. In cases where either of these models fails, our approach may not function optimally. However, because our refinement modules are lightweight, simple, and model-agnostic, incorporating more advanced methods for coarse alignment could further enhance performance. Moreover, while existing pose-free view synthesis methods (Chen & Lee, 2023; Smith et al., 2023; Hong et al., 2024; Ye et al., 2024) share a limitation that requires camera intrinsic parameters, exploring joint optimization of both camera extrinsic and intrinsic is a promising direction. Finally, our method is not designed for dynamic scenes, but we believe exploring feed-forward dynamic scene reconstruction is a promising direction.
@@ A.6 Limitations and Future Work
Figure 7: Pose estimation qualitative results.
Figure 8: Qualitative results on RealEstate-10K with 6 input views.
Figure 9: Qualitative results on RealEstate-10K with 12 input views.